OpenStudy (anonymous):

x^2 - 10^x + 16 = -5 Can someone write this out in standard form and show their work. I had trouble learning this during the year and would like to be prepared.

OpenStudy (anonymous):

I am pretty sure all you have to do to put that in standard form is to get one side to equal 0 ...so just add 5 to both sides of the equation

OpenStudy (anonymous):

Thank you so Paynes!

OpenStudy (wolf1728):

x² -10x +21 = 0 by quadratic formula a = 1 b=-10 c=21 x = [-(-10) ±sqrt(100 - 4 * 1 * 21) / (2*1)] x = [10 ±sqrt(16)/2] x = [10 +4 / 2] = 7 x = [10 -4 / 2] = 3